If a scientist has a 94 mL solution of 0.1 M HCl, then to a good approximation, what will be the pH of the solution? Enter your answer to the nearest hundredth.
Answer #1

Assumption: given solution contains only HCl or non-affecting constituents in the solution.

HCl(aq) -----> H+(aq) + Cl-(aq)

HCl is a strong acid and hence it will dissociate completly.

we have 94 mL = 94*10^-3 L

of 0.1 M HCl, then total number of moles = 94*10^-3 L * 0.1 moles/L = 94*10^-4 moles

Thus total number of moles of H+ = 94*10^-4 moles.

So, Molarity of H+ [H+] = 94*10^-4 moles / 94*10^-3 L = 0.10 moles/L = 0.10 M

As per the defination of pH = -log[H+] = -log(0.10) = 1.00

Thus the pH of solution is 1.00